OpenStudy (anonymous):

Solve. 3x^2 – 48 = 0 a. x = 3 and x = 16 b. x = 4 and x = –4 c. x = 16 and x = 4 d. x = 4 and x = 3

OpenStudy (anonymous):

\[3x^2 – 48 = 0 \]\[3x^2 = 48\]\[x^2= 48/3\]\[x^2=16\]\[x=\pm \sqrt{16}\]\[x=\pm 4\] So the answer is b.

razor99 (razor99):

its corect 100%

OpenStudy (anonymous):

thank u, i understand now